OpenStudy (katie0146722):

Which shows the fractions in order from least to greatest? five-twelfths, two-ninths, five-sixths two-ninths, five-twelfths, five-sixths five-twelfths, two-ninths, five-sixths five-sixths, two-ninths, five-twelfths five-sixths, five-twelfths, two-ninths

OpenStudy (anonymous):

First off, We need to find a common denominator

OpenStudy (katie0146722):

five-twelfths, two-ninths, five-sixths

OpenStudy (anonymous):

Yes but not simplified enough

OpenStudy (katie0146722):

let me try again

OpenStudy (katie0146722):



OpenStudy (katie0146722):

chinese boy can u please hlp? I will give u medal

OpenStudy (anonymous):

Okay!

OpenStudy (katie0146722):

thnx

OpenStudy (anonymous):

Wait me a minute please.

OpenStudy (anonymous):

aw so unfair! lmao

OpenStudy (anonymous):

Now I come.

OpenStudy (katie0146722):

I will give u a medal soon girl

OpenStudy (katie0146722):

there @GirlByte

OpenStudy (anonymous):

Aw thanks so much :D

OpenStudy (katie0146722):

hihihi

OpenStudy (anonymous):

5/12 2/9 5/6 30/72 16/72 50/72

OpenStudy (anonymous):

so 2/9 is smallest, 5/12 is middle, 5/6 is highest

OpenStudy (anonymous):

Dang katie, you are going rambo giving me medals!

OpenStudy (katie0146722):

hihihi yw

OpenStudy (katie0146722):

thats none of the answer choices

OpenStudy (katie0146722):

five-twelfths, two-ninths, five-sixths two-ninths, five-twelfths, five-sixths five-twelfths, two-ninths, five-sixths five-sixths, two-ninths, five-twelfths five-sixths, five-twelfths, two-ninths

OpenStudy (anonymous):

dudeso 2/9 is smallest, 5/12 is middle, 5/6 is highest

OpenStudy (anonymous):

I already answered it, it's the second.

OpenStudy (katie0146722):

thnx so so so so so much @trumpetmaster7777

OpenStudy (katie0146722):

and thnx @Chineseboy15 4 trying

OpenStudy (anonymous):

np

OpenStudy (katie0146722):

can u guys help me on 1 more? please

OpenStudy (katie0146722):

Compare: eight and eleven-fifteenths to eight and thirteen-twentieths less than greater than equal to None of the above

OpenStudy (anonymous):

Oh, sorry, I made some mistakes. The right answer is \[2/9<5/12<5/6\]

OpenStudy (katie0146722):

thnx

OpenStudy (anonymous):

HI KATIE

OpenStudy (anonymous):

The two fractions are \[8\frac{ 11 }{ 15 }, 8\frac{ 13 }{ 20 }\] Now to make them be \[\frac{ 131 }{ 15 }, \frac{ 173 }{ 20 }\] Then\[\frac{ 131\times4 } { 15\times4 }, \frac{ 173\times3 }{ 20\times3 }\] \[\frac{ 524 }{ 60 } >\frac{ 519 }{ 60 }\]

OpenStudy (katie0146722):

hi lol

OpenStudy (katie0146722):

thank u